OpenStudy (angela210793):

I know this is a veeeeeeeeery stupid question but plss don't comment just answer: c-(c-2)/3=? for c=-1 is 0 right?

OpenStudy (amistre64):

\[\frac{c-(c-2)}{3}=?; c=-1\]

OpenStudy (angela210793):

C-2 C- ------=? when C=-1 this is it 3

OpenStudy (anonymous):

yup

OpenStudy (angela210793):

kk thnx ^_^